OpenStudy (anonymous):

Find the Solution of 2cos^2x + 3sin x =0

OpenStudy (anonymous):

2cos^2x + 3sin x =0 2(1-sin^2x) + 3sinx = 0 -2sin^2x + 3sinx +2 = 0 2sin^2x -3sinx -2 = 0 (2sinx +1)(sinx-2) = 0 Solve for x

OpenStudy (anonymous):

sinx - 2 =0 sinx = 2

OpenStudy (anonymous):

2sin^2 x = 1 sin^2 x = 1/2

OpenStudy (anonymous):

2sin^2 x = 1 <- not really!

OpenStudy (anonymous):

sinx = 2 not posiible

OpenStudy (anonymous):

sin^2 x =- 1/2

OpenStudy (anonymous):

sin x = -1/2

OpenStudy (anonymous):

x = sin (180 +150) and sin(360-330)

OpenStudy (anonymous):

x=arcsin(-1/2)

OpenStudy (anonymous):

sin^2 x =- 1/2 <- incorrect sin x = -1/2 <- correct x = sin (180 +150) and sin(360-330) <- not correct It should be \(x = sin^{-1} (-1/2)\)

OpenStudy (anonymous):

or x =11pie/6

OpenStudy (anonymous):

I made it correct u see

OpenStudy (anonymous):

we will have two values for x

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

When you get the step sinx = -1/2 You should take arc sine for both sides to solve x sin^-1 (sinx) = sin^-1(-1/2) x = -30 or x = 180+30

OpenStudy (anonymous):

This is because there are two values for x in which sinx = -1/2

OpenStudy (anonymous):

wat r they!!

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

we need to find solution such npie + (-1)^n y

OpenStudy (anonymous):

What is sin^-1 (-1/2)?

OpenStudy (anonymous):

i dont knw

OpenStudy (anonymous):

What value of x can give you -1/2 when you take sine of that value?

OpenStudy (anonymous):

No idea!

OpenStudy (anonymous):

What given you 1/2 when you take sine of that value??

OpenStudy (anonymous):

30

OpenStudy (anonymous):

pie/6

OpenStudy (anonymous):

And??

OpenStudy (anonymous):

wat??

OpenStudy (anonymous):

sin(360-330) = sin330 = sin11pie/6

OpenStudy (anonymous):

sin(360-330) ≠ sin330 !!!

OpenStudy (anonymous):

sin(2pie - 11pie/6) =?

OpenStudy (anonymous):

sin(2pie - 11pie/6) = sin (pi/6) = 1/2

OpenStudy (anonymous):

ok......PLzz teach me hw to take inverse function

OpenStudy (anonymous):

Think in this way, when you take sine of a function, what values would give you -1/2?

OpenStudy (anonymous):

ok...then it should lie in 3rd or 4th quadrant

OpenStudy (anonymous):

Yes, but are the values?

OpenStudy (anonymous):

values hw to find..

OpenStudy (anonymous):

Consider quadrant I, sin (pi/6) = 1/2 Consider quadrant II, sin (pi - pi/6) = 1/2 Consider quadrant III, sin (pi + pi/6) = -1/2 Consider quadrant IV, sin (2pi - pi/6) = -1/2 Got it?

OpenStudy (anonymous):

thxxx a lot

OpenStudy (anonymous):

You're welcome.